# Convergence of the open WASEP stationary measure without Liggett’s condition

### v1

@@ 3.1. Convergence of One-point Distributions
This completes the proof of tightness for 1.3. It is important to note that, while Corwin and Knizel state Lemma 5.1 for the scaling conditions in 1.4, the proof relies on an attractive coupling of ASEP and a sandwiching argument for the diffusively scaled random walks $h_{u,-v,w,r}^{(N)}(x)$ and $h_{-v,v,w,r}^{(N)}(x)$ which is unchanged when we allow $w$ and $r$ to vary.
Finally, we can prove that our sequence of measures converges weakly to $H_{u,v}(\cdot)-H_{u,v}(0)$, completing the proof of 1.3.
Proof of 1.3.
We must show that, when $u+v>0$ and $w,r>0$, the finite dimensional distributions of $\{h_{u,v,w,r}^{(N)}(\cdot)-h_{u,v,w,r}^{(N)}(0)\}$ converge in distribution to those of $H_{u,v}(\cdot)-H_{u,v}(0)$. Using a standard result in the convergence of continuous stochastic processes (see, for example, Karatzas and Shreve [KS91], Ch. 2 Theorem 4.15), convergence in finite dimensional distributions with tightness of the sequence, from 1.3 $(1)$, implies that $\{h_{u,v,w,r}^{(N)}(\cdot)-h_{u,v,w,r}^{(N)}(0)\}\implies H_{u,v}(\cdot)-H_{u,v}(0)$ on $(C[0,1],\mathcal{B}(C[0,1])).$
For any $d\in\mathbb{Z}_{\geq 1}$ and $0<x_{1}<...<x_{d}\leq 1$ and $x^{(N)}=N^{-1}\lfloor Nx\rfloor$, we use $\mathbb{P}^{(N)}$ to denote the law of the vector $\left(h_{u,v,w,r}^{(N)}(x_{1}^{(N)})-h_{u,v,w,r}^{(N)}(0),...,h_{u,v,w,r}^{(N)}(x_{d}^{(N)})-h_{u,v,w,r}^{(N)}(0)\right)$. Consequently, using 1.3 $(1)$, we know that every subsequence of this collection has a convergent sub-subsequence. We know that the corresponding sequence of Laplace transforms converges to a unique limit, which is the Laplace transform of the open KPZ stationary measure on an open interval (3.1). Therefore, applying 2.6, we see that when $u+v>0$ and $w,r>0$, $\{h^{(N)}_{u,v,w,r}(\cdot)-h^{(N)}_{u,v,w,r}(0)\}\implies H_{u,v}(\cdot)-H_{u,v}(0).$

### v2

@@ 4. Convergence of the Laplace Transform
This completes the proof of tightness for 1.1. It is important to note that, while [CK24] states Lemma 5.1 for the scaling conditions in 1.2, the proof relies on an attractive coupling of ASEP and a sandwiching argument for the diffusively scaled random walks $h_{u,-v,w,r}^{(N)}(x)$ and $h_{-v,v,w,r}^{(N)}(x)$ which is unchanged when we allow $w$ and $r$ to vary.
Finally, we can prove that our sequence of measures converges weakly to $H_{u,v}(\cdot)-H_{u,v}(0)$, completing the proof of 1.1.
Proof of 1.1.
We must show that, when $u+v>0$ and $w,r>0$, the finite dimensional distributions of $\{h_{u,v,w,r}^{(N)}(\cdot)-h_{u,v,w,r}^{(N)}(0)\}$ converge in distribution to those of $H_{u,v}(\cdot)-H_{u,v}(0)$. Using a standard result in the convergence of continuous stochastic processes (see, for example, Karatzas and Shreve [KS91], Ch. 2 Theorem 4.15), convergence in finite dimensional distributions with tightness of the sequence, from 1.1 $(1)$, implies that $\{h_{u,v,w,r}^{(N)}(\cdot)-h_{u,v,w,r}^{(N)}(0)\}\implies H_{u,v}(\cdot)-H_{u,v}(0)$ on $(D[0,1],\mathcal{B}(D[0,1])).$
For any $d\in\mathbb{Z}_{\geq 1}$ and $0<x_{1}<...<x_{d}\leq 1$ and $x^{(N)}=N^{-1}\lfloor Nx\rfloor$, we use $\mathbb{P}^{(N)}$ to denote the law of the vector $\left(h_{u,v,w,r}^{(N)}(x_{1}^{(N)})-h_{u,v,w,r}^{(N)}(0),...,h_{u,v,w,r}^{(N)}(x_{d}^{(N)})-h_{u,v,w,r}^{(N)}(0)\right)$. Consequently, using 1.1 $(1)$, we know that every subsequence of this collection has a convergent sub-subsequence. We know that the corresponding sequence of Laplace transforms converges to a unique limit, which is the Laplace transform of the open KPZ stationary measure on an open interval (3.1). Therefore, applying 2.4, we see that when $u+v>0$ and $w,r>0$, $\{h^{(N)}_{u,v,w,r}(\cdot)-h^{(N)}_{u,v,w,r}(0)\}\implies H_{u,v}(\cdot)-H_{u,v}(0).$